If One number is 7 less than anoher. If 4 times the smaller number plus 2 times the larger number is 62, find the 2 numbers. will it be 8,15 did I do this right?
Answer by rmromero(383)


What is asked in the problem?
find the two number
Given:
one number is 7 less than another.
4 times the smaller number plus 2 times the larger number is 62.
Representation:
Let x = the larger number
x - 7 = the smaller number
Write and Equation: Translate the sentence into mathematical equation.
4 times the smaller number plus 2 times the larger number is 62.
4(x - 7) + 2(x) = 62
4x - 28 + 2x = 62
6x = 90
x = 15 ----------->> Larger number
x - 7 = the smaller number, x = 15
15 - 7 = 8 ---------->> Smaller number
Checking :

4(8) + 2(15) = 62
32 + 30 = 62
62 = 62 -------->> True


Therefore, your answer is correct.